“Blast it!” said David Wilson, president of Teledex Company. “We’ve just lost the bid on the Koopers job by $4,000. It seems we’re either too high to get the job or too low to make any money on half the jobs we bid.”

Teledex Company manufactures products to customers’ specifications and uses a job-order costing system. The company uses a plantwide predetermined overhead rate based on direct labor cost to apply its manufacturing overhead (assumed to be all fixed) to jobs. The following estimates were made at the beginning of the year:

Department Fabricating Machining Assembly Total Plant
Manufacturing Overhead 378,000 432,000 97,200 907,200
Direct Labor 216,000 108,000 324,000 648,000

Jobs require varying amounts of work in the three departments. The Koopers job, for example, would have required manufacturing costs in the three departments as follows:

Department Fabricating Machining Assembly Total Plant
Direct Materials 4,600 200 3,000 7,800
Direct Labor 6,000 500 7,800 14,300
Manufacturing Overhead ? ? ? ?

Required:

1. Using the company's plantwide approach:

a. Compute the plantwide predetermined rate for the current year.

b. Determine the amount of manufacturing overhead cost that would have been applied to the Koopers job.

2. Suppose that instead of using a plantwide predetermined overhead rate, the company had used departmental predetermined overhead rates based on direct labor cost. Under these conditions:

a. Compute the predetermined overhead rate for each department for the current year.

b. Determine the amount of manufacturing overhead cost that would have been applied to the Koopers job.

4. Assume that it is customary in the industry to bid jobs at 150% of total manufacturing cost (direct materials, direct labor, and applied overhead).

a. What was the company’s bid price on the Koopers job using a plantwide predetermined overhead rate?

b. What would the bid price have been if departmental predetermined overhead rates had been used to apply overhead cost?
Answer #1

1a)usind plantwide approach ,Predetermined overhead rate =Estimated overhead /estimated direct labor cost

                                                      = 907200/648000

                                                     = 1.4 or 140% of direct labor cost

b)Overhead Applied =Actual labor cost *predetermined overhead rate

                      = 14300 *140%.

                      = 20020

2a)

Fabricating Machining Assembly`
Manufacturing overhead 378000 432000 97200
Direct labor cost 216000 108000 324000
Departmental overhead rate 378000/216000=1.75 or 175% 432000/108000=4 or 400% .30 or 30%

b)

Fabricating Machining Assembly` Total
Actual direct labor cost 6000 500 7800 14300
Overhead applied 6000*175%= 10500 500*400%= 2000 7800*30%= 2340 14840

4)

a)

Direct material 7800
Direct labor 14300
overhead applied 20020
Total cost 42120
Bid price 42120*150%= 63180

b)

Direct material 7800
Direct labor 14300
overhead applied 14840
Total cost 36940
Bid price 36940*150%= 55410
Answer #2

1.a)

Plantwide Predetermined Rate = Budgeted Plantwide Cost/Budgeted Total Labor cost

= $907,200/$648,000

= 1.4/Direct Labor

1.b)

Manufacturing Overhead Applied to Koopers Job = Actual Direct Labor cost Incurred * predetermined rate.

= $14,300 * 1.4

= $20,020

2.a)

Predetermined Overhead Rate for each department = Budgeted overhead cost/Budgeted Labor Cost

For Fabrication Department:-

= $378,000/$216,000

= 1.75

For Machining Department:-

= $432,000/$108,000

= 4

For Assembly Department:-

= $97,200/$324,000

= .3

2.b)

Manufacturing overhead Applied to Koopers Job:-

Fabrication Department = $6,000*1.75 =$10,500

Machining Department = $500*4 = $ 2,000

Assembly Department = 7,800*.3 = $ 2,340

Total Overhead Applied = $14,840

4.a)

Company's Bid price on Koopers Job Under plantwide predetermined rate:-

Direct Materials = $7,800

Direct Labor = $14,300

Overhead Applied = $20,020 (calculated in 1.b)

Total Cost = $42,120

Bid price of Company = $42,120+($42,120*150%)   

= $105,300

4.b)

Company's Bid price for Koopers Job under Departmental predetermined rate:-

  Direct Material = $7,800

Direct Labor = $14,300

Overhead Applied = $14,840 (calculated in 2.b)

Total cost = $36,940

Bid price of Company = $36,940+($36,940*150%)

= $92,350